# La RoSA: Enhancing LLM Efficiency via Layerwise Rotated Sparse Activation

### v1

@@ Abstract
Activation sparsity can reduce the computational overhead and memory transfers during the forward pass of Large Language Model (LLM) inference. Existing methods face limitations, either demanding time-consuming recovery training that hinders real-world adoption, or relying on empirical magnitude-based pruning, which causes fluctuating sparsity and unstable inference speed-up. This paper introduces LaRoSA (Layerwise Rotated Sparse Activation), a novel method for activation sparsification designed to improve LLM efficiency without requiring additional training or magnitude-based pruning. We leverage layerwise orthogonal rotations to transform input activations into rotated forms that are more suitable for sparsification. By employing a Top-K selection approach within the rotated activations, we achieve consistent model-level sparsity and reliable wall-clock time speed-up. LaRoSA is effective across various sizes and types of LLMs, demonstrating minimal performance degradation and robust inference acceleration. Specifically, for LLaMA2-7B at 40% sparsity, LaRoSA achieves a mere 0.17 perplexity gap with a consistent 1.30× wall-clock time speed-up, and reduces the accuracy gap in zero-shot tasks compared to the dense model to just 0.54%, while surpassing TEAL by 1.77% and CATS by 17.14%.
@@ 1 Introduction
Large Language Models (LLMs) have achieved significant advancements across various real-world machine learning applications and leaderboards (DeepSeek-AI, 2025; OpenAI, 2024; Team, 2024; Qwen et al., 2025). However, billions of parameters in these models pose significant challenges for efficient LLM inference and serving, especially in terms of memory footprint and computational overhead. Recent efforts on boosting LLM inference have predominantly focused on low-bit model quantization (Frantar et al., 2022; Lin et al., 2024b; Ashkboos et al., 2024b; Liu et al., 2024b; Lin et al., 2024a) and model weights pruning methods (Frantar & Alistarh, 2023; Ashkboos et al., 2024a; Ma et al., 2023; Fang et al., 2024). Such techniques enhance inference speed and save memory usage by either employing low-precision parameter formats or reducing the total number of parameters. Exploiting activation sparsity (Kurtz et al., 2020; Li et al., 2022; Chen et al., 2023) is another promising approach to improve inference efficiency. Since channels in the weight matrix corresponding to zero-valued activations are not utilized during the forward pass, LLM inference can be accelerated by: (1) avoiding transfer such parameters between memory and computation units, and (2) omitting the matrix computations associated with those weights channels. These speed-ups could be achieved by utilizing sparse patterns that are compatible with hardware.
@@
Recent studies (Mirzadeh et al., 2024; Liu et al., 2023; Zhang et al., 2024) have observed a high degree of activation sparsity in the ReLU-based Transformers. Based on this, Dejavu (Liu et al., 2023) trains a predictor for each layer to preemptively identify weight channels corresponding to zero-valued activations.  (Mirzadeh et al., 2024) induce high sparsity by replacing non-ReLU activations with ReLU. However, these methods do not achieve full sparsity in the models and require substantial additional training to recover model performance. Some other works have turned their attention to magnitude-based pruning. CATS (Lee et al., 2024) utilizes empirical distribution histogram to obtain the cutoff threshold in MLP blocks, and zero out activations below that threshold. TEAL (Liu et al., 2024a) further extends this concept to all input activations in both Attention and MLP blocks. However, offline cutoff thresholds are heavily related to calibration tokens, and lack accuracy and consistence in generating optimal sparse patterns for unseen tokens to achieve a target sparsity level. Moreover, inconsistent activation sparsity can lead to unstable inference speed-ups, reducing efficiency in LLM inference. Driven by these limitations, we raise the following question: Is it possible to quickly, effectively, and consistently leverage activation sparsity to enhance the efficiency of LLMs without additional training or empirical thresholds?
In this work, we present LaRoSA (Layerwise Rotated Sparse Activation), a training-free activation sparsification method that produces optimal sparse patterns and consistent generation speed-up for LLM inference. As depicted in Figure 1, our method uses an orthogonal rotation matrix ${\mathbf{Q}}$ to transform the input activation, enabling more effective activation sparsification. We apply PCA to obtain the rotation matrix ${\mathbf{Q}}$ for each layer and employ a Top-K function to acquire stable activation sparsity. A residual adapter is added to realize layer-specific independent rotations. In the end, the rotation matrix ${\mathbf{Q}}$ can be absorbed by the corresponding weight matrix, eliminating the additional computational overhead introduced by ${\mathbf{Q}}$. Our main contributions are:
@@ 2 Related Work
Activation Sparsity Activation sparsity refers to the phenomenon where a considerable portion of a model’s hidden states are zero-valued. This property naturally emerges in the intermediate states of ReLU-based MLP blocks (Li et al., 2022). Liu et al. (2023) utilized activation sparsity to accelerate LLM decoding speed by omitting the transfer of weight channels associated with zero-valued entries to GPU registers. Furthermore, Song et al. (2023) and Alizadeh et al. (2024) leveraged activation sparsity to facilitate CPU offloading, significantly reducing memory transfer overhead between CPUs and GPUs. However, it is noteworthy that modern leading LLMs (Team, 2024; Qwen et al., 2025) utilize more complex nonlinear functions (e.g., SwiGLU (Shazeer, 2020)) which do not inherently produce sparsity (Mirzadeh et al., 2024). Consequently, most of the ReLU-based activation sparsification methods are not applicable to these models.
Inducing Sparsity in non-ReLU Models Mirzadeh et al. (2024) replaced SwiGLU with ReLU in LLMs, followed by continued pre-training to induce activation sparsity. Song et al. (2025) and Song et al. (2024) introduced techniques such as activation regularization to further enhance sparsity in adapted models. Zhang et al. (2024) experimented with various activation functions and identified Squared ReLU (So et al., 2021) as the most effective alternative. CATS (Lee et al., 2024) realizes a desired sparsity level in SwiGLU based MLP output activations through empirical distribution and magnitude thresholding. TEAL (Liu et al., 2024a) further extends this concept to the input activations of self-attention and MLP blocks, and prunes low-magnitude activations. Wang et al. (2024) combines magnitude pruning and quantized activations and presents scaling laws for sparsely activated LLMs. However, these methods either rely on substantial post-training and calibration, or produce unstable model-level sparsity, resulting in sparse activated LLMs that lack efficiency and flexibility.
Rotation-based Model Compression Recent works have focused on using orthogonal rotation matrices to mitigate performance drops in LLM pruning and quantization. SliceGPT (Ashkboos et al., 2024a) multiplies each weight matrix by an orthogonal matrix to reduce the embedding dimension, and use a residual adapter to achieve computational invariance. QuaRot (Ashkboos et al., 2024b) use random Hadamard rotation to redistribute outliers across all channels. SpinQuant (Liu et al., 2024b) suggests training the orthogonal matrix instead of using a random Hadamard matrix to further enhance quantizability.
@@ 3.1 Activation Sparsity in LLMs
In this work, we focus on the Transformer architecture LLMs which are composed of repeated decoder layers (Touvron et al., 2023). In subsequent sections, unless otherwise specified, a ‘layer’ refers to the decoder layer that repeatedly appears after the token embedding layer, and a ‘block’ refers to an Attention block or a MLP block from a decoder layer. We assume that the network is constructed using a pre-norm approach, meaning that a LayerNorm or RMSNorm operation precedes each block. For clarity, ${\bm{h}}_{1}$ and ${\bm{h}}_{2}$ refer to the Attention block input activations, and ${\bm{h}}_{3}$ and ${\bm{h}}_{4}$ to the MLP block input activations. Figure 2 illustrates the construction of such blocks and layers. Typically, the forward pass in each block is performed by linear mappings between input tensors and weight matrices, which can be expressed as:
@@
where ${\mathbf{X}}\in\mathbb{R}^{N\times D_{\text{in}}}$ denote the input tensor of $N$ tokens, ${\mathbf{W}}\in\mathbb{R}^{D_{\text{in}}\times D_{\text{out}}}$ is the weight matrix, and ${\mathbf{Y}}\in\mathbb{R}^{N\times D_{\text{out}}}$ is the output tensor. $D_{\text{in}}$ and $D_{\text{out}}$ are hidden sizes of input and output tensor. The matrix multiplication shown in Equation 1 is identical for each $\mathbf{W}_{\text{q}}$, $\mathbf{W}_{\text{k}}$, $\mathbf{W}_{\text{v}}$, $\mathbf{W}_{\text{o}}$ in Attention blocks and $\mathbf{W}_{\text{up}}$, $\mathbf{W}_{\text{gate}}$, $\mathbf{W}_{\text{down}}$ in MLP blocks. Let ${\bm{x}}\in\mathbb{R}^{D_{\text{in}}}$ denotes the input activation of a single token from ${\mathbf{X}}$. The activation sparsity $p\in[0,1]$ of ${\bm{x}}$ is defined as the proportion of zero-valued elements:
@@
where $x_{i}$ is the i-th element of ${\bm{x}}$, and the indicator function $\mathbf{1}({\bm{x}}_{i}=0)$ outputs 1 if the activation value is equal to zero and 0 otherwise. As highlighted in  (Mirzadeh et al., 2024), non-ReLU based LLMs inherently exhibit extremely low activation sparsity. Recent studies (Lee et al., 2024; Liu et al., 2024a) have assumed that low-magnitude activations exert minimal effects on model outputs and have shown that magnitude-based pruning is empirically effective. Typically, the activation sparsity ins these methods is defined as:
@@ 3.2 Computational Invariance Transformation
The computational invariance theorem, as described in (Ashkboos et al., 2024a) Theorem 1, asserts that the activations between Transformer blocks (i.e., Attention and MLP) can be transformed using an orthogonal matrix ${\mathbf{Q}}$ without altering the model’s output. This theorem remains valid even when RMSNorm is applied between the blocks. The reason is that RMSNorm divides activations by their norms, and a orthogonal transformation ${\mathbf{Q}}$ does not affect these norms, since ${\mathbf{Q}}{\mathbf{Q}}^{\top}={\mathbf{Q}}^{\top}{\mathbf{Q}}={\mathbf{I}}$ and $\|{\bm{x}}{\mathbf{Q}}\|=\sqrt{{\bm{x}}{\mathbf{Q}}{\mathbf{Q}}^{\top}{\bm{x}}^{\top}}=\sqrt{{\bm{x}}{\bm{x}}^{\top}}=\|{\bm{x}}\|$. Therefore, we have the the commutation property in RMSNorm operation:
@@
In forward pass, we can merge the additional transformation ${\mathbf{Q}}$ into weight matrices to avoid the extra computation induced by ${\mathbf{Q}}$. As depicted in Figure 2, if $\mathbf{W}_{\text{o}}$ is right-multiplied by an orthogonal matrix $\mathbf{Q}$, this will make the Attention block produces a rotated activation ${\mathbf{X}}\mathbf{Q}$ as output. We can cancel this effect by left-multiplying the weight matrices of subsequent MLP blocks (i.e., $\mathbf{W}_{\text{up}}$, $\mathbf{W}_{\text{gate}}$) by $\mathbf{Q}^{\top}$, so the output changes induced by $\mathbf{Q}$ are neutralized, and the original computation results in each block remain unchanged. This process is same for $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{q}}$, $\mathbf{W}_{\text{k}}$, $\mathbf{W}_{\text{v}}$.
@@ 4.1 Motivation: Rethinking Low-Magnitude Pruning
We conduct experiments on LLaMA3-8B using magnitude pruning methods, using 16 sequences of 512 tokens from WikiText2 (Merity et al., 2016) for various sparsity settings. We identify several key observations and visualizing them in Figure 3.
Ambiguity and Inaccuracy in Defining Low Magnitude. The definition of low magnitude influences how activations are sparsified. As shown in Figure 3(a), cutoff thresholds obtained from offline calibration fluctuate significantly as the layer goes deeper. The thresholds for Attention and MLP blocks within the same layer also exhibit great variation across different sparsity levels. Using these inaccurate thresholds can lead to severe mismatches with target sparsity. Figure 3(b) compares cutoff thresholds for activation pruning using TEAL and TopK methods at various sparsity levels. TEAL uses a fixed threshold from offline calibration, shown as a dashed line, while TopK computes thresholds during testing, shown as scatter points. It’s evident that offline calibrated thresholds hardly align with the actual needed thresholds. In most cases, tokens actually require higher or lower thresholds than the calibrated ones. At sequence beginnings, actual thresholds are extremely lower than calibrated ones, explaining why methods like TEAL struggle to sparsify initial tokens of a sequence during prefill stage of LLM inference.
@@
Magnitude and Channel Importance Misconception. Magnitude pruning methods (Liu et al., 2024a; Lee et al., 2024) typically assume that low-magnitude activations have a minor impact on model outputs, leading to the practice of zeroing out entries that fall below a certain threshold. However, as proved in Wanda (Sun et al., 2023) and SliceGPT (Ashkboos et al., 2024a), pruning solely based on the magnitude of weights or activations is not optimal. Low-magnitude activations can still substantially affect the output if their corresponding channels in the weight matrix have large norms. We further demonstrate in Appendix A that empirical output errors from magnitude pruning methods are consistently higher than theoretical output errors.

### v2

@@ Abstract
Activation sparsity can reduce the computational overhead and memory transfers during the forward pass of Large Language Model (LLM) inference. Existing methods face limitations, either demanding time-consuming recovery training that hinders real-world adoption, or relying on empirical magnitude-based pruning, which causes fluctuating sparsity and unstable inference speed-up. This paper introduces LaRoSA (Layerwise Rotated Sparse Activation), a novel method for activation sparsification designed to improve LLM efficiency without requiring additional training or magnitude-based pruning. We leverage layerwise orthogonal rotations to transform input activations into rotated forms that are more suitable for sparsification. By employing a Top-K selection approach within the rotated activations, we achieve consistent model-level sparsity and reliable wall-clock time speed-up. LaRoSA is effective across various sizes and types of LLMs, demonstrating minimal performance degradation and robust inference acceleration. Specifically, for LLaMA2-7B at 40% sparsity, LaRoSA achieves a mere 0.17 perplexity gap with a consistent 1.30× wall-clock time speed-up, and reduces the accuracy gap in zero-shot tasks compared to the dense model to just 0.54%, while surpassing TEAL by 1.77% and CATS by 17.14%. Source code: https://github.com/alibaba/EfficientAI.
@@ 1 Introduction
Large Language Models (LLMs) have achieved significant advancements across various real-world machine learning applications and leaderboards (deepseekr1; openai2024gpt4technicalreport; llama3; qwen2). However, billions of parameters in these models pose significant challenges for efficient LLM inference and serving, especially in terms of memory footprint and computational overhead. Recent efforts on boosting LLM inference have predominantly focused on low-bit model quantization (frantar-gptq; lin2023awq; ashkboos2024quarot; liu2024spinquant; lin2024duquant) and model weights pruning methods (frantar-sparsegpt; ashkboos2024slicegpt; ma2023llmpruner; fang2024maskllm). Such techniques enhance inference speed and save memory usage by either employing low-precision parameter formats or reducing the total number of parameters. Exploiting activation sparsity (pmlr-v119-FATReLU; Li2022TheLN; Chen2023SparseViT) is another promising approach to improve inference efficiency. Since channels in the weight matrix corresponding to zero-valued activations are not utilized during the forward pass, LLM inference can be accelerated by: (1) avoiding transfer such parameters between memory and computation units, and (2) omitting the matrix computations associated with those weights channels. These speed-ups could be achieved by utilizing sparse patterns that are compatible with hardware.
@@
Recent studies (reluSB; dejavu; Zhang2024ReLU2WD) have observed a high degree of activation sparsity in the ReLU-based Transformers. Based on this, Dejavu (dejavu) trains a predictor for each layer to preemptively identify weight channels corresponding to zero-valued activations.  (reluSB) induce high sparsity by replacing non-ReLU activations with ReLU. However, these methods do not achieve full sparsity in the models and require substantial additional training to recover model performance. Some other works have turned their attention to magnitude-based pruning. CATS (lee2024cats) utilizes empirical distribution histogram to obtain the cutoff threshold in MLP blocks, and zero out activations below that threshold. TEAL (liu2024teal) further extends this concept to all input activations in both Attention and MLP blocks. However, offline cutoff thresholds are heavily related to calibration tokens, and lack accuracy and consistence in generating optimal sparse patterns for unseen tokens to achieve a target sparsity level. Moreover, inconsistent activation sparsity can lead to unstable inference speed-ups, reducing efficiency in LLM inference. Driven by these limitations, we raise the following question: Is it possible to quickly, effectively, and consistently leverage activation sparsity to enhance the efficiency of LLMs without additional training or empirical thresholds?
In this work, we present LaRoSA (Layerwise Rotated Sparse Activation), a training-free activation sparsification method that produces optimal sparse patterns and consistent generation speed-up for LLM inference. As depicted in Figure˜1, our method uses an orthogonal rotation matrix $\mathbf{Q}$} to transform the input activation, enabling more effective activation sparsification. We apply PCA to obtain the rotation matrix $\mathbf{Q}$} for each layer and employ a Top-K function to acquire stable activation sparsity. A residual adapter is added to realize layer-specific independent rotations. In the end, the rotation matrix $\mathbf{Q}$} can be absorbed by the corresponding weight matrix, eliminating the additional computational overhead introduced by $\mathbf{Q}$}. Our main contributions are:
@@ 2 Related Work
Activation Sparsity Activation sparsity refers to the phenomenon where a considerable portion of a model’s hidden states are zero-valued. This property naturally emerges in the intermediate states of ReLU-based MLP blocks (Li2022TheLN). dejavu utilized activation sparsity to accelerate LLM decoding speed by omitting the transfer of weight channels associated with zero-valued entries to GPU registers. Furthermore, song2023powerinfer and alizadeh2024llm_in_a_flash leveraged activation sparsity to facilitate CPU offloading, significantly reducing memory transfer overhead between CPUs and GPUs. However, it is noteworthy that modern leading LLMs (llama3; qwen2) utilize more complex nonlinear functions (e.g., SwiGLU (shazeer2020SwiGLU)) which do not inherently produce sparsity (reluSB). Consequently, most of the ReLU-based activation sparsification methods are not applicable to these models.
Inducing Sparsity in non-ReLU Models reluSB replaced SwiGLU with ReLU in LLMs, followed by continued pre-training to induce activation sparsity. song2025prosparse and song2024turbosparse introduced techniques such as activation regularization to further enhance sparsity in adapted models. Zhang2024ReLU2WD experimented with various activation functions and identified Squared ReLU (squared_relu) as the most effective alternative. CATS (lee2024cats) realizes a desired sparsity level in SwiGLU based MLP output activations through empirical distribution and magnitude thresholding. TEAL (liu2024teal) further extends this concept to the input activations of self-attention and MLP blocks, and prunes low-magnitude activations. wang2024qsparses combines magnitude pruning and quantized activations and presents scaling laws for sparsely activated LLMs. However, these methods either rely on substantial post-training and calibration, or produce unstable model-level sparsity, resulting in sparse activated LLMs that lack efficiency and flexibility.
Rotation-based Model Compression Recent works have focused on using orthogonal rotation matrices to mitigate performance drops in LLM pruning and quantization. SliceGPT (ashkboos2024slicegpt) multiplies each weight matrix by an orthogonal matrix to reduce the embedding dimension, and use a residual adapter to achieve computational invariance. QuaRot (ashkboos2024quarot) use random Hadamard rotation to redistribute outliers across all channels. SpinQuant (liu2024spinquant) suggests training the orthogonal matrix instead of using a random Hadamard matrix to further enhance quantizability.
@@ 3.1 Activation Sparsity in LLMs
In this work, we focus on the Transformer architecture LLMs which are composed of repeated decoder layers (llama2). In subsequent sections, unless otherwise specified, a ‘layer’ refers to the decoder layer that repeatedly appears after the token embedding layer, and a ‘block’ refers to an Attention block or a MLP block from a decoder layer. We assume that the network is constructed using a pre-norm approach, meaning that a LayerNorm or RMSNorm operation precedes each block. For clarity, ${\bm{h}}_{1}$ and ${\bm{h}}_{2}$ refer to the Attention block input activations, and ${\bm{h}}_{3}$ and ${\bm{h}}_{4}$ to the MLP block input activations. Figure˜2 illustrates the construction of such blocks and layers. Typically, the forward pass in each block is performed by linear mappings between input tensors and weight matrices, which can be expressed as:
@@
where ${\mathbf{X}}\in\mathbb{R}^{N\times D_{\text{in}}}$ denote the input tensor of $N$ tokens, ${\mathbf{W}}\in\mathbb{R}^{D_{\text{in}}\times D_{\text{out}}}$ is the weight matrix, and ${\mathbf{Y}}\in\mathbb{R}^{N\times D_{\text{out}}}$ is the output tensor. $D_{\text{in}}$ and $D_{\text{out}}$ are hidden sizes of input and output tensor. The matrix multiplication shown in Equation˜1 is identical for each $\mathbf{W}_{\text{q}}$, $\mathbf{W}_{\text{k}}$, $\mathbf{W}_{\text{v}}$, $\mathbf{W}_{\text{o}}$ in Attention blocks and $\mathbf{W}_{\text{up}}$, $\mathbf{W}_{\text{gate}}$, $\mathbf{W}_{\text{down}}$ in MLP blocks. Let ${\bm{x}}\in\mathbb{R}^{D_{\text{in}}}$ denotes the input activation of a single token from $\mathbf{X}$}. The activation sparsity $p\in[0,1]$ of $\bm{x}$} is defined as the proportion of zero-valued elements:
@@
where $x_{i}$ is the i-th element of $\bm{x}$}, and the indicator function $\mathbf{1}({\bm{x}}_{i}=0)$ outputs 1 if the activation value is equal to zero and 0 otherwise. As highlighted in  (reluSB), non-ReLU based LLMs inherently exhibit extremely low activation sparsity. Recent studies (lee2024cats; liu2024teal) have assumed that low-magnitude activations exert minimal effects on model outputs and have shown that magnitude-based pruning is empirically effective. Typically, the activation sparsity ins these methods is defined as:
@@ 3.2 Computational Invariance Transformation
The computational invariance theorem, as described in (ashkboos2024slicegpt) Theorem 1, asserts that the activations between Transformer blocks (i.e., Attention and MLP) can be transformed using an orthogonal matrix $\mathbf{Q}$} without altering the model’s output. This theorem remains valid even when RMSNorm is applied between the blocks. The reason is that RMSNorm divides activations by their norms, and a orthogonal transformation $\mathbf{Q}$} does not affect these norms, since ${\mathbf{Q}}{\mathbf{Q}}^{\top}={\mathbf{Q}}^{\top}{\mathbf{Q}}={\mathbf{I}}$ and $\|{\bm{x}}{\mathbf{Q}}\|=\sqrt{{\bm{x}}{\mathbf{Q}}{\mathbf{Q}}^{\top}{\bm{x}}^{\top}}=\sqrt{{\bm{x}}{\bm{x}}^{\top}}=\|{\bm{x}}\|$. Therefore, we have the the commutation property in RMSNorm operation:
@@
In forward pass, we can merge the additional transformation $\mathbf{Q}$} into weight matrices to avoid the extra computation induced by $\mathbf{Q}$}. As depicted in Figure˜2, if $\mathbf{W}_{\text{o}}$ is right-multiplied by an orthogonal matrix $\mathbf{Q}$, this will make the Attention block produces a rotated activation ${\mathbf{X}}\mathbf{Q}$ as output. We can cancel this effect by left-multiplying the weight matrices of subsequent MLP blocks (i.e., $\mathbf{W}_{\text{up}}$, $\mathbf{W}_{\text{gate}}$) by $\mathbf{Q}^{\top}$, so the output changes induced by $\mathbf{Q}$ are neutralized, and the original computation results in each block remain unchanged. This process is same for $\mathbf{W}_{\text{down}}$ and $\mathbf{W}_{\text{q}}$, $\mathbf{W}_{\text{k}}$, $\mathbf{W}_{\text{v}}$.
@@ 4.1 Motivation: Rethinking Low-Magnitude Pruning
We conduct experiments on LLaMA3-8B using magnitude pruning methods, using 16 sequences of 512 tokens from WikiText2 (wikitext) for various sparsity settings. We identify several key observations and visualizing them in Figure˜3.
Ambiguity and Inaccuracy in Defining Low Magnitude. The definition of low magnitude influences how activations are sparsified. As shown in Figure˜3(a), cutoff thresholds obtained from offline calibration fluctuate significantly as the layer goes deeper. The thresholds for Attention and MLP blocks within the same layer also exhibit great variation across different sparsity levels. Using these inaccurate thresholds can lead to severe mismatches with target sparsity. Figure˜3(b) compares cutoff thresholds for activation pruning using TEAL and TopK methods at various sparsity levels. TEAL uses a fixed threshold from offline calibration, shown as a dashed line, while TopK computes thresholds during testing, shown as scatter points. It’s evident that offline calibrated thresholds hardly align with the actual needed thresholds. In most cases, tokens actually require higher or lower thresholds than the calibrated ones. At sequence beginnings, actual thresholds are extremely lower than calibrated ones, explaining why methods like TEAL struggle to sparsify initial tokens of a sequence during prefill stage of LLM inference.
@@
Magnitude and Channel Importance Misconception. Magnitude pruning methods (liu2024teal; lee2024cats) typically assume that low-magnitude activations have a minor impact on model outputs, leading to the practice of zeroing out entries that fall below a certain threshold. However, as proved in Wanda (wanda) and SliceGPT (ashkboos2024slicegpt), pruning solely based on the magnitude of weights or activations is not optimal. Low-magnitude activations can still substantially affect the output if their corresponding channels in the weight matrix have large norms. We further demonstrate in Appendix˜A that empirical output errors from magnitude pruning methods are consistently higher than theoretical output errors.